In the adjoining circuit, the switch, which had been closed for a sufficiently long time for steady state to be reached, is opened at time t = 0. Determine the following, as a function of time:

(a) The current I L(t) through the inductor, and (b) The voltage v R(t) across the 1k Ohm resistor.

I=0 Rs=5.12 [26) Vs= + Ro= 1 k 2 20 V 1 H 0000 vr(t)
Answer #1

solerioni Giren circuit is an Rg=52 t=0 following: ato 31926) + Đ V=20volt TH vele Ro= { V (H) At t=0-(moment just before swiThe curent (t) can be be expressed as, Pe (t) = f (as) - {12 (20) - Plotzet where, 2= time constant of the cincerit= Roo 10-3